OpenStudy (anonymous):

In an arithmetic sequence consisting of 20 numbers is the first number 13 and the differential 7. a)Determine the first five number b)Determine the twelfth number c) Determine the total number sequence

OpenStudy (akashdeepdeb):

n = 20 [no. of terms] d = 7 [common difference] a = 13 [first term] Now the Sequence would look like: 13, 20, 27 ...... till 20 terms. You can come up with a formula for finding the nth term. A_n = a + (n-1)d [It is a formula for the nth element, do you want the explanation to the formula?] A_n = 13 + (n-1)*7 So first term would be A_1 = 13 + (1-1)*7 = 13 A_2 = 13 + (2-1)*7 = 20 ... find all terms like this.. :D Understood? :)

OpenStudy (nirmalnema):

a+(n-1)d a=13 d=7 a1=13 a2=13+7=20 a3=20+7=27 a4=27+7=34 a5=34+7=41

OpenStudy (nirmalnema):

for n=12 13+(12-1)7 13+77 =90

OpenStudy (anonymous):

and the sum is 146 right??

OpenStudy (nirmalnema):

sum of what

OpenStudy (akashdeepdeb):

@nirmalnema You are not allowed to give all answers on this website, explain it. @sarapetrini The sum of all terms is given by the formula in arithmetic sequence S_n = n*{2a+(n-1)*d]/2 You have all the numbers, calculate! :D

OpenStudy (nirmalnema):

ok.... @AkashdeepDeb

OpenStudy (anonymous):

i need the answer because i suck on mathematics, im on a test right now, im not gonna use this mathematics ever! im gonna be a spanish teacher so im never gonna use it again, i just need to pass this horrible test :S lol.. thank you both!!! :)))))

OpenStudy (anonymous):

@phi @preetha since I cant seem to access my mod account, see if you want to deal with this stuff.

OpenStudy (anonymous):

Akash, of course nirma just went and gave a direct answer to her next question, so lieing is another characteristic of this user.

OpenStudy (akashdeepdeb):

Yeah? Hopefully, some suggestions I posted are going to change that. :)